a simple supported beam of 8m pan carries the total load of 23kn/m includingof of ot self weight .use m30 and fe 415 design the beam .
1
Expert's answer
2021-07-29T22:21:01-0400

Self-weight of footing

"P = 23 kN/m*8m = 184 kN\\
P_T= 1.2 *184 kN = 220.8 kN\\"

From the M30 and the Fe415 design the beam

"W_T=\frac{P_T}{A}\\
120=\frac{220.8}{A} \implies A = 1.84 \\"

"Area = S^2 \implies (B)= \sqrt{1.84}= 1.35646"

The size chosen can be "1.5 m \times 1.5 m"

Therefore area "= 2.25m^2"

Pressure

"P= FA \\
P =496.8 N/m^2"